You are an analyst employed by an airplane manufacturer that last year sold 40,000 ATR-72

aircrafts at $100,000 each. Your market research indicates that:

I. the price elasticity of demand for your aircrafts in −0.5. (or +0.5 in absolute value);

II. the income elasticity of demand for your aircrafts is +3.7; and

III. the cross price elasticity for your aircrafts with respect to the price of a comparable jet

manufactured by a competitor is +1.6.

A. Suppose that you expect a ceteris paribus decrease in average incomes of 10% this

year compared to last year. How many aircrafts do you estimate that your company will

sell this year? How will it impact total revenues? 6 marks

B. Assume now that you do not think incomes will change, but that you expect your

competitor will decrease his price by 4%. Assuming that your company does not change

the price of its aircrafts, how many would you expect your company will sell this year?


1
Expert's answer
2021-11-29T19:10:07-0500

A.

"E(ID)=\frac{\Delta Q}{\Delta I}"


"-37=\frac{Q2-Q2}{Q2+Q1}"


"3.7=\frac{\Delta Q}{-10}"

"\Delta Q=-37"


"-37=\frac{Q2-Q2}{Q2+Q1}=\frac{Q2-40000}{Q2+40000}"

-38Q2=-1520 000

Q2=40000


B.

"E(AB)d=\frac{\Delta QA}{\Delta QB}"

"1.6=\frac{\Delta QA}{-4}"

"\Delta QA=-6.4"



"-6.4=\frac{Q2-Q2}{Q2+Q1}=\frac{Q2-40000}{Q2+40000}"

-7.4Q2=-296000

Q2=40000